OpenStudy (anonymous):

f(x) = (x-15)^5 If f(x)=f(g(x)), how do I find f(x) and g(x) ?

OpenStudy (anonymous):

g(x) = x-15 f(x) = x^5. You can see it by looking at the function f(x) as a nested function, consisting of g(x) and f(x).